OpenStudy (anonymous):

Solve the system by substitution. Check your solution. a-1.2b=-3 0.2b+0.6A=12 a. (15, 15) c. (13, 12) b. (10, 12) d. (7, 9)

OpenStudy (mathmale):

Dear NicoleAnn: Your system, a-1.2b=-3 0.2b+0.6A=12 could be made easier to solve by eliminating all the decimal fractions. To accomplish this, multiply both equations by 10. Then: 10a - 12b = -30 6a + 2b = 120 One of the easier ways to solve this would be to multiply the 2nd eq'n by 6, so that the b terms will be -12b and +12b respectively. Adding together the 2 equations will eliminate b and make it easy to solve for a. Good luck!

OpenStudy (anonymous):

@mathmale Thank you! I got 15,15, Is this corect?

OpenStudy (mathmale):

Let's look at the process of checking your "solution" of a problem such as this one. You're wondering whether (15,15) is a solution. Choose either one of the equations you originally posted here. For example, take a-1.2b=-3 and substitute a=15 and b=15. Is the resulting equation true or false? If true, you have a solution; if false, something's wrong. You could check your sol'n in the other equation in exactly the same way.

OpenStudy (mathmale):

If we have a-1.2b=-3 0.2b+0.6A=12 I'd do the same thing as before: Multiply each equation by 10: 10a-12b=-30 6 a +2b = 120 Next, I'd multiply the 2nd equation by 6: 36a + 12 b = 720 and add the first equation to it: 36a + 12 b = 720 10a-12b=-30 The resulting equation is 46a=690, so a = 15. Now we need to find b: Substitute a = 15 into the 2nd equation: 10(15) -12b =-30 Solve the resulting equation for b. Then check your "solution" again in the manner discussed above. Suppose we get (15,15) again: Substitute a=15 and b=15 into 36a + 12 b = 720. Is the equation then correct? 36(15) + 12(15) = 720 ??

OpenStudy (mathmale):

So, in fewer words :) Yes, your answer is correct!!